OpenStudy (anonymous):

solve a/5 - 18 = 2 a divided by 5 minus 18 = 2

OpenStudy (anonymous):

a = 2(5-18) a = 10-36 a=-26

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

a/5 - 18 = 2 ===> a/5=20 ===> a=100